Metallic iron contains approximately 1029 atoms per cubic meter. The magnetic moment
of each iron atom is 1.8 × 10−23 Am−1
. If all the dipoles were perfectly aligned what
would be the magnetisation M and the magnetic moment of a bar 10 cm long and 1cm2
in
cross-section?
1
Expert's answer
2019-03-25T08:47:50-0400

Magnetization is the magnetic moment of a unit volume, so that M = 1029*1.8*10-23 = 1.8*106 (A/m), and the magnetic moment of the bar is Mb = MV = 1.8*106*0.1*10-4 = 18 (Am2).